OpenStudy (anonymous):

Write an equation for line L in point intercept form L is parallel to y=2x L=(4,-2)

OpenStudy (anonymous):

L is parallel to y=2x therefore, slope = 2 and L=(4,-2) y-(-2) =(slope)(x-4) y+2=2(x-4) y+2=2x-8 y=2x-8-2 = 2x-10